# Expert System for Bitcoin Forecasting: Integrating Global Liquidity via TimeXer Transformers

### v1

@@ Abstract
Bitcoin price forecasting is characterized by extreme volatility and non-stationarity, making it hard for univariate time-series prediction models over long horizons. By integrating Global M2 Liquidity—aggregated from 18 major economies as exogenous variable we improve the predictions over long horizons.
Using TimeXer, we compare a liquidity-conditioned univariate forecasting model (TimeXer-Exog) against state-of-the-art univariate models, LSTM, N-BEATS, PatchTST, and standard TimeXer. Experiments conducted on daily Bitcoin price data from January 2020 to August 2025 demonstrate that explicit macroeconomic conditioning significantly stabilizes long-horizon forecasts.
At a 70-day forecast horizon, the proposed TimeXer-Exog model achieves a Mean Squared Error (MSE) of $10.814\times 10^{7}$ which outperformed the univariate TimeXer baseline by over 89%. The results show that explicitly conditioning deep learning models on Global Liquidity as an exogenous variable gives substantial improvements in long-horizon Bitcoin price forecasting.
Keywords: Bitcoin forecasting, Global liquidity, TimeXer, Transformers, Deep learning
@@ 1 Introduction
Early attempts to forecast Bitcoin prices were done using statistical tools, especially ARIMA models (Sagheer2025; Xu2025). ARIMA is good at picking up short-term linear patterns, like the ones in a relatively calm equity index over a few weeks. However, bitcoin does not behave like that. Its price is noisy and very volatile, which statistical models and tools fail to capture. ARIMA requires stationarity and linear structure, this is exactly what bitcoin isn’t. Large rallies, sudden crashes, or regime shifts often get smoothed away (Pratas2023; Mizdrakovic2024). As a result, these models can perform well over a short horizon, their accuracy drops once forecast horizon is about 60 or 90 days (Mousa2025; Kareem2025).
@@
Researchers moved RNN models, which were designed to handle sequences. LSTM and GRU models, were successful in learning the non-linear patterns from the data (Bagheri2025; Xu2025). They can, for example, react differently to a slow upward trend than to a sudden spike triggered by a regulatory headline. These models are not perfect. Because they process data step by step, they scale poorly and struggle when dependencies stretch far back in time. Capturing relationships across hundreds of days still remains hard, and unreliable (Zhao2022; Zheng2025).
@@
More recently, attention has shifted toward Transformer-based models, time series forecasting moved to self-attention rather than strict sequence processing (Nie2022; Zhao2022). Instead of remembering the past one step at a time, these models can look across an entire window at once. PatchTST takes this idea further by breaking the time series into patches, small segments that preserve local structure while keeping computation manageable (Nie2022). For time series forecasting which require long look-back windows, Transformer models have repeatedly outperformed LSTM-based approaches, especially for long-horizon predictions (Nie2022; Zheng2025).
@@
Long-term forecasting improves when models are allowed to look beyond price history itself. The TimeXer framework reflects this shift. It combines patch-wise attention for the main series with cross-attention for external variables, making it possible to integrate data that arrive at different frequencies and are not neatly aligned in time (Wang2024). This is closer to how real financial systems operate. Macroeconomic indicators do not update daily, yet markets respond to them gradually and sometimes unevenly.
@@
Global liquidity, often measured by M2 money supply, has attracted particular attention (Gu2025). The logic is intuitive, even if the relationship is not perfectly stable. Empirical work using time-varying Granger causality suggests that the influence of M2 on Bitcoin prices comes and goes, strengthening during expansionary regimes and fading at other times (Gu2025). Cointegration analyses go a step further, pointing to a long-run equilibrium relationship in which Bitcoin prices respond more than proportionally to liquidity growth (Kokabian2025). Importantly, these effects do not appear overnight. The impact of M2 unfolds with lags, which fits the idea that liquidity works its way through portfolios and risk appetite before reaching crypto markets (Kokabian2025).
We wanted a forecasting approach that combines long-horizon modeling with explicit macroeconomic context. In this paper we integrate global M2 money supply as an exogenous variable within a Transformer-based framework. Using TimeXer and global liquidity as exogenous variable, we can have better bitcoin price predictions over long horizons.
From a broader financial engineering perspective, this direction reflects a quiet change in how intelligent systems are defined. Accuracy matters, but so does context. Analysts look at bitcoin prices often in relation to liquidity cycles, central bank policy, and broader risk sentiment. This is what differentiates a macroeconomist from a black-box predictor. Moving from a univariate setup to a multivariate expert framework, and by allowing the model to explicitly attend to global monetary conditions, the proposed approach aims to deliver insights that are more robust to the drift and structural change that plague many traditional forecasting tools.
@@ 2.1 Bitcoin forecasting using ARIMA
Bitcoin price predictions was initially done using ARIMA, partly because it was already a familiar tool in financial time series analysis (Kareem2025; Xu2025). For short-term movements, especially when prices drift without too much drama, ARIMA can be very reliable. One can imagine fitting it over a calm stretch of daily log returns and getting something that looks sensible for the next few days. The trouble starts when the market becomes volatile. Bitcoin prices are rarely stationary for too long, sharp jumps or sudden collapses violate the linear structure ARIMA assumes (Sagheer2025; Pratas2023). During volatile episodes, the model has a habit of smoothing away extreme moves, as if the crash or rally were just noise rather than the main event (Kareem2025). Forecasts also lose credibility quickly as the forecast horizon increases, which makes ARIMA a weak candidate for longer-term planning (Mousa2025).
@@ 2.2 Bitcoin forecasting using CNN, RNN, and LSTM models
As attention shifted toward more flexible approaches, deep learning models began to dominate the discussion (Bagheri2025). RNNs, and LSTMs in particular, seemed well suited to the task. They were designed to remember patterns over longer periods of time, and in practice they handled Bitcoin’s nonlinear behavior better than classical models (Hochreiter1997; Mardjo2024). Some studies paired CNNs with LSTMs, letting convolutional layers pick up local price patterns, such as short bursts of momentum, before passing them to a recurrent structure (Sagheer2025; Bagheri2025). Interestingly, several comparative analyses found that GRUs often matched or even outperformed LSTMs, while using fewer parameters and training more efficiently (Kaur2025; Mohammadjafari2024; Nafisah2025). These models share a common weakness. Because they process sequences step by step, they become cumbersome when dependencies stretch far back in time, which is often the case in crypto markets which are shaped by long liquidity cycles.
@@ 2.3 Bitcoin forecasting using Transformer-based models
Transformer models entered the scene as a response to these limitations (Zheng2025). By relying on self-attention rather than strict sequential processing, they can examine long windows of data in parallel. This makes it easier to capture distant dependencies without the training bottlenecks typical of RNNs. In volatile settings like cryptocurrency markets, time-series Transformers have delivered encouraging results (Zheng2025). In PatchTST we divide the input series into patches, a practical compromise that preserves local structure while keeping attention costs under control (Nie2022). Across a range of long-horizon tasks, transformer models have repeatedly outperformed LSTM baselines, producing forecasts that are not only more accurate but also less erratic over time (Nie2022). That said, architecture alone does not explain everything, and context still matters.
@@ 2.4 Incorporating exogenous variables in Bitcoin forecasting
Long-term Bitcoin forecasting rarely works if only price history is considered. Researchers have experimented with on-chain activity, sentiment measures, and macroeconomic indicators to reflect broader market forces (Mahfooz2024; Grubisic2025). Among these, global liquidity, typically proxied by M2 money supply, stands out (Gu2025). The connection is not mechanical, but there is evidence of a long-run equilibrium relationship between M2 and Bitcoin prices, with the influence becoming more pronounced during periods of monetary expansion (Gu2025; Kokabian2025). Importantly, these effects tend to emerge with delays rather than instant reactions. Transformer based models like TimeXer make it easier to incorporate such variables by allowing cross-attention between price dynamics and external drivers (Wang2024). This line of work points toward forecasting systems that treat Bitcoin not as an isolated series, but as part of a wider monetary environment, where global liquidity an play a role in shaping long-horizon price behavior.
@@ 3.1 The TimeXer Architecture
TimeXer was built with a fairly practical goal in mind: making it easier to bring external information into time series forecasting without tearing apart the Transformer framework. TimeXer reuses the familiar Transformer backbone and adapts how inputs are represented and attended to. The result feels less like a radical redesign and more like a careful rethinking of how different data streams should talk to each other.
@@
As shown in Figure 1, the model processes the input through two parallel pathways before combining them in the Transformer backbone. The endogenous (a) divides the Bitcoin price series into non-overlapping patches and adds a randomly initialized global token, forming the main temporal representation. At the same time, the exogenous (b) processes the Global Liquidity series separately, embedding it into variate-wise tokens that retain information about the external macroeconomic driver.
@@ 3.2 Baseline Models
To benchmark performance, several established deep learning were used for univariate forecasting:
@@
  • PatchTST: This Transformer variant utilizes subseries-level patches along with channel independence to enhance long-term forecasting accuracy (Nie2022).
  • LSTM: A foundational recurrent architecture designed to overcome the vanishing gradient problem and efficiently capture long-term dependencies in sequential data (Hochreiter1997; Mahfooz2024).
  • N-BEATS: A purely deep learning architecture consisting of a very deep stack of fully-connected layers linked by residual connections (Oreshkin2019).
Classical linear models such as ARIMAX are known to suffer from stationary constraints and recursive forecast instability in highly volatile and non-stationary financial series; they are included only as diagnostic baselines and are not expected to compete with attention-based models at extended forecast horizons.(melina2024comparative)
@@ 3.3 Bitcoin and Global Liquidity Relationship
The integration of Global M2 Liquidity is justified by its documented long-run influence on Bitcoin pricing. A cointegration analysis established a strong relationship, characterized by a long-run elasticity estimate of 2.65, meaning a 1% increase in the M2 money supply is associated with a 2.65% increase in the price of Bitcoin (Kokabian2025). The stability of this long-run relationship is validated by a statistically significant Error Correction Term (ECT), $\lambda^{\prime}=-0.12$ (Kokabian2025). The magnitude of this coefficient demonstrates that approximately 12% of any deviation from the long-run equilibrium is corrected monthly.
@@
Figure 2 highlights a clear lead-lag relationship between Global Liquidity and Bitcoin’s price action. Shifting the M2 liquidity data back by 12 weeks, the alignment of major cyclical peaks and troughs can clearly be observed. This observation suggests that changes in the global liquidity acts as a leading indicator, with a transmission delay of approximately one fiscal quarter before the liquidity is fully reflected in Bitcoin’s price. TimeXer leverages this 12-week lead as a predictive ”look-ahead” window, thus improving the predictions.
@@ 3.4 Statistical Evaluation: Model Confidence Set (MCS)
To evaluate the predictive performance of the proposed TimeXer-Exog model against the set of benchmark models ($\mathcal{M}_{0}=\{\text{TimeXer, PatchTST, LSTM, N-BEATS}\}$), we use Model Confidence Set (MCS) Hansen2011MCS.
Unlike traditional tests (e.g., the Diebold-Mariano test (Diebold1995)), which can suffer from the multiple comparisons problem and data snooping bias when many models are evaluated (White2000; Sullivan1999), the Model Confidence Set (MCS) procedure determines a subset of models $\mathcal{M}^{*}\subset\mathcal{M}_{0}$ that contains the ”true best” models with a specified confidence level $(1-\alpha)$. We iteratively eliminate models found to be significantly inferior until the null hypothesis of Equal Predictive Ability (EPA) cannot be rejected for the remaining set (Hansen2011MCS).
Since we use long horizon forecasting (Horizon $h=70$), the forecast errors are expected to exhibit serial correlation up to lag h1h-1. To account for this dependency structure and prevent spurious significance, we utilize the Block Bootstrap method (Kunsch1989) within the MCS framework. We set the block size equal to the forecast horizon ($h$) to preserve the internal correlation structure of the error series during resampling (Politis2004). We set the significance level $\alpha=0.10$. This means models with a p-value <0.10<0.10 are excluded from the Superior Set Model (SSM).
@@ 4.1 Data Description
To ensure strict out-of-sample evaluation, the final 120 daily observations, from January 22, 2025, to May 21, 2025, were reserved exclusively for testing. All models were trained only on historical data prior to this window.
@@
The exogenous variable employed in this study is Global Liquidity, constructed as an aggregate measure of broad money supply across a diversified set of major and emerging economies. Specifically, the indicator includes M2 (or the closest available broad money proxy) from the following economies: the United States, Euro Area, China, Japan, United Kingdom, Canada, Australia, India, South Korea, Brazil, Russia, Switzerland, Mexico, Indonesia, Turkey, Saudi Arabia, Argentina, and South Africa.

### v2

@@ Abstract
Bitcoin price forecasting is characterized by extreme volatility and non-stationarity, making it hard for univariate time-series prediction models over long-horizons. By integrating Global M2 Liquidity as an exogenous variable, aggregated from 18 major economies, we improve the predictions over long-horizons.
Using TimeXer, we compare a liquidity-conditioned univariate forecasting model (TimeXer-Exog) against state-of-the-art univariate models, Long Short-Term Memory(LSTM), Neural Basis Expansion Analysis for interpretable Time-Series (N-BEATS), Patch Time-Series Transformer(PatchTST), and standard TimeXer. Experiments conducted on daily Bitcoin price data from January 2020 to August 2025 demonstrate that explicit macroeconomic conditioning significantly stabilizes long-horizon forecasts.
At a 70-day forecast horizon, the proposed TimeXer-Exog model achieves a Mean Squared Error (MSE) of $10.814$, scaled by $10^{7}$, which outperformed the univariate TimeXer baseline by over 89%. The results show that explicitly conditioning deep learning models on Global Liquidity as an exogenous variable gives substantial improvements in long-horizon Bitcoin price forecasting.
Keywords: Bitcoin forecasting, Global liquidity, TimeXer, Transformers, Deep learning, Macroeconomic conditioning, Long-horizon forecasting
@@ 1 Introduction
Early attempts to forecast Bitcoin prices were made using statistical tools, especially the ARIMA models (Sagheer et al., 2025; Xu, 2025). ARIMA is good at picking up short-term linear patterns, like the ones in a relatively calm equity index over a few weeks. However, bitcoin does not behave like that. Its price is noisy and very volatile, which statistical models and tools do not capture. ARIMA requires stationarity and linear structure, which is exactly what bitcoin lacks. Large rallies, sudden crashes, or regime shifts often get smoothed away (Pratas et al., 2023; Mizdrakovic, 2024). As a result, these models can perform well over a short horizon, their accuracy drops once the forecast horizon is about 60 or 90 days (Mousa, 2025; Kareem and Aue, 2025).
@@
Researchers moved to RNN models, which were designed to handle sequences. Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) models were successful in learning the non-linear patterns from the data (Bagheri and Giudici, 2025; Xu, 2025). They can, for example, react differently to a slow upward trend than to a sudden spike triggered by a regulatory headline. These models are not perfect. Because they process data step by step, they scale poorly and struggle when dependencies stretch far back in time. Capturing relationships across hundreds of days still remains hard, and unreliable (Zhao et al., 2022; Zheng, 2025).
@@
More recently, attention has shifted toward Transformer-based models, time series forecasting moved to self-attention rather than strict sequence processing (Nie, 2022; Zhao et al., 2022). Instead of remembering the past one step at a time, these models can look across an entire window at once. PatchTST takes this idea further by breaking the time series into patches, small segments that preserve local structure while keeping computation manageable (Nie, 2022). For time series forecasting which require long look-back windows, Transformer models have repeatedly outperformed LSTM-based approaches, especially for long-horizon predictions (Nie, 2022; Zheng, 2025).
@@
Long-term forecasting improves when models are allowed to look beyond price history itself. The TimeXer framework reflects this shift. It combines patch-wise attention for the main series with cross-attention for external variables, making it possible to integrate data that arrive at different frequencies and are not neatly aligned in time (Wang, 2024). This is closer to how real financial systems operate. Macroeconomic indicators do not update on a daily basis, yet markets tend to respond to them gradually and unevenly.
@@
Global liquidity is often measured by M2 money supply, which has particularly attracted attention (Gu and Chen, 2025). Although the relationship is not perfectly stable, the logic is intuitive. According to empirical work using time-varying Granger causality, the influence of M2 on Bitcoin prices is not constant. It strengthens during expansionary regimes and weakens in other regimes (Gu and Chen, 2025). Co-integration analyses go further by pointing towards a long-run equilibrium relationship in which Bitcoin prices respond more than proportionally to liquidity growth (Kokabian, 2025). It is important to note that these effects do not come into the picture suddenly. The impact of M2 unfolds with lags which corresponds to the idea that before reaching the crypto market, liquidity works its way through portfolios and risk appetite (Kokabian, 2025).
Our aim is to develop an expert forecasting system that transcends simple pattern recognition by integrating long-horizon modeling with explicit macroeconomic context. By embedding Global M2 Liquidity as an exogenous variable within the TimeXer framework, we enable the model to emulate the decision-making process of a human macro-strategist. Much like an expert analyst who weighs monetary policy shifts against technical price action, our system uses cross-attention mechanisms to dynamically attend to global liquidity cycles, thereby stabilizing predictions over longer horizons.
From a broader financial engineering perspective, this represents a meaningful evolution from univariate forecasting to intelligent decision support. Context is as critical as computation. While black-box predictors often fail during structural breaks, this approach incorporates domain-specific expert knowledge—specifically, the lead-lag relationship between money supply and risk assets. By moving from a univariate setup to a multivariate expert framework, the proposed approach delivers insights that are robust to the drift and structural changes common in digital asset markets, effectively functioning as an automated expert for capital allocation.
@@ 2.1 Bitcoin forecasting using ARIMA
Bitcoin price predictions were conducted using ARIMA initially, largely because it was already a familiar tool in financial time series analysis (Kareem and Aue, 2025; Xu, 2025). For short-term movements, especially when prices drift without too much drama, ARIMA can be very reliable. One can imagine fitting it over a calm stretch of daily log returns and getting something that looks sensible for the next few days. The trouble starts when the market becomes volatile. Bitcoin prices are rarely stationary for too long, sharp jumps or sudden collapses violate the linear structure ARIMA assumes (Sagheer et al., 2025; Pratas et al., 2023). During volatile episodes, the model has a habit of smoothing away extreme moves, as if the crash or rally were just noise rather than the main event (Kareem and Aue, 2025). Forecasts also lose credibility quickly as the forecast horizon increases, which makes ARIMA a weak candidate for longer-term planning (Mousa, 2025).
@@ 2.2 Bitcoin forecasting using CNN, RNN, and LSTM models
As attention shifted toward more flexible approaches, deep learning models began to dominate the discussion.(Bagheri and Giudici, 2025). Models such as RNNs and LSTMs performed particularly well and seemed to suit the task. They remembered patterns over longer periods of time and handled Bitcoin’s nonlinear behavior better than classical models (Hochreiter and Schmidhuber, 1997; Mardjo and Choksuchat, 2024). Researchers combined CNNs with LSTMs, letting convolutional layers pick up local price patterns, such as short bursts of momentum, before passing them to a recurrent structure (Sagheer et al., 2025; Bagheri and Giudici, 2025). Several comparative analyses found that GRUs often matched or even outperformed LSTMs, while using fewer parameters and training more efficiently (Kaur, 2025; Mohammadjafari, 2024; Nafisah, 2025). However, these models have a common drawback. Since they process sequences step by step, they often become cumbersome when dependencies are from the past, which is often the case in crypto markets because they are shaped by long liquidity cycles.
@@ 2.3 Bitcoin forecasting using Transformer-based models
To tackle these limitations, Transformer models entered the scene (Zheng, 2025). They relied on self-attention rather than strict sequential processing and they examined longer windows of data in parallel. This makes it more efficient to capture distant dependencies without the trouble of training bottlenecks typical of RNNs. Time-series Transformers have delivered encouraging and positive results in cryptocurrency markets which have volatile settings (Zheng, 2025). In PatchTST, we divide the input series into patches. It can be considered as a practical compromise that preserves local structure while keeping attention costs under control (Nie, 2022). Results across a range of long-horizon tasks indicate that Transformer models have repeatedly outperformed LSTM baselines, producing forecasts that are not only more accurate but also less erratic over time (Nie, 2022). In conclusion, architecture alone does not explain everything, and context still matters.
@@ 2.4 Incorporating exogenous variables in Bitcoin forecasting
Relying solely on price data is not effective for long-term Bitcoin predictions. Researchers have experimented with on-chain activity, sentiment measures, and macroeconomic indicators to reflect broader market forces (Mahfooz and Phillips, 2024; Grubišić, 2025). Global liquidity turned out to have a significant linkage (Gu and Chen, 2025). The connection is not mechanical, but there is evidence of a long-run equilibrium relationship between M2 and Bitcoin prices, with the influence becoming more pronounced during periods of monetary expansion (Gu and Chen, 2025; Kokabian, 2025). It is important to note that these effects tend to emerge with lags rather than instant reactions. Transformer-based models like TimeXer make it easier to incorporate such variables by allowing cross-attention between price dynamics and external drivers (Wang, 2024). This line of work points toward forecasting systems that treat Bitcoin not as an isolated series, but as part of a wider monetary environment, where global liquidity plays a role in shaping long-horizon price behavior.
@@ 3.1 The TimeXer Architecture
TimeXer was built to ease bringing external information into time series forecasting without disturbing the Transformer framework. TimeXer reuses the familiar Transformer backbone and adapts how inputs are represented and attended to. The result feels less like a radical redesign and more like a careful rethinking of how different data streams should talk to each other.
@@
As shown in Figure 1, the model processes the input through two parallel pathways before combining them in the Transformer backbone. The endogenous (a) divides the Bitcoin price series into non-overlapping patches and adds a randomly initialized global token, forming the main temporal representation. In parallel, the exogenous (b) processes the Global Liquidity series separately, embedding it into variate-wise tokens that retain information about the external macroeconomic driver.
@@ 3.2 Baseline Models
To benchmark performance, several established deep learning models were used for univariate forecasting:
@@
  • PatchTST: This Transformer variant utilizes subseries-level patches along with channel independence to enhance long-term forecasting accuracy (Nie, 2022).
  • LSTM: A recurrent architecture designed to mitigate the vanishing gradient problem and capture long-term dependencies in sequential data (Hochreiter and Schmidhuber, 1997; Mahfooz and Phillips, 2024).
  • N-BEATS: A purely deep learning architecture consisting of a very deep stack of fully-connected layers linked by residual connections (Oreshkin et al., 2019).
Exclusion of Linear Baselines.
Classical linear models, such as ARIMA and ARIMAX, were excluded from the final comparative analysis for the long-horizon tasks ($H=70$). These models rely on recursive (iterative) forecasting strategies to generate multi-step predictions. In high-volatility environments like cryptocurrency markets, recursive strategies suffer from severe error accumulation, where slight deviations in early steps propagate and amplify over the forecast horizon (Hyndman and Athanasopoulos, 2018). Furthermore, the stationarity transformations required for ARIMA to function (e.g., differencing) often result in a loss of the long-term trend information that is critical for the specific task of tracking global liquidity cycles. Consequently, we focus our benchmarking exclusively on deep learning architectures capable of direct multi-step forecasting.
@@ 3.3 Bitcoin and Global Liquidity Relationship
The integration of Global M2 Liquidity is justified by its documented long-run influence on Bitcoin pricing. A cointegration analysis established a strong relationship, characterized by a long-run elasticity estimate of 2.65, meaning a 1% increase in the M2 is associated with a 2.65% increase in the price of Bitcoin (Kokabian, 2025). The stability of this long-run relationship is validated by a statistically significant Error Correction Term (ECT), $\lambda^{\prime}=-0.12$ (Kokabian, 2025). The magnitude of this coefficient demonstrates that approximately 12% of any deviation from the long-run equilibrium is corrected monthly.
@@
Figure 2 highlights a clear lead-lag relationship between Global Liquidity and Bitcoin’s price action. Shifting the M2 liquidity data back by 12 weeks, the alignment of the major cyclical peaks and troughs can be clearly observed. This observation suggests that changes in the global liquidity act as a leading indicator, with a transmission delay of approximately one fiscal quarter before the liquidity is fully reflected in Bitcoin’s price. TimeXer leverages this 12-week lead as a predictive ”look-ahead” window, thus improving the predictions.
@@ 3.4 Statistical Evaluation: Model Confidence Set (MCS)
To evaluate the predictive performance of the proposed TimeXer-Exog model against the set of benchmark models ($\mathcal{M}_{0}=\{\text{TimeXer, PatchTST, LSTM, N-BEATS}\}$), we use MCS Hansen et al., 2011.
Unlike traditional tests (e.g., the Diebold-Mariano test (Diebold and Mariano, 1995)), which can suffer from the multiple comparisons problem and data snooping bias when many models are evaluated (White, 2000; Sullivan et al., 1999), the MCS procedure determines a subset of models $\mathcal{M}^{*}\subset\mathcal{M}_{0}$ that contains the ”true best” models with a specified confidence level $(1-\alpha)$. We iteratively eliminate models found to be significantly inferior until the null hypothesis of Equal Predictive Ability (EPA) cannot be rejected for the remaining set (Hansen et al., 2011).
Since we use long-horizon forecasting (Horizon $h=70$), the forecast errors are expected to exhibit serial correlation up to lag h1h-1. We implemented the Block Bootstrap method (Künsch, 1989) within the MCS framework to account for this dependency structure and prevent spurious significance. In order to preserve the internal correlation structure of the error series during resampling (Politis and White, 2004), we set the block size equal to the forecast horizon ($h$) and the significance level to $\alpha=0.10$. This means models with a p-value <0.10<0.10 are excluded from the Superior Set Model (SSM).
@@ 4.1 Data Description
To ensure strict out-of-sample evaluation, the dataset was partitioned into training, validation, and test sets. The final 120 daily observations, from January 22, 2025, to May 21, 2025, were reserved exclusively for testing. The preceding 120 days, from September 24, 2024, to January 21, 2025, were used as the validation set for hyperparameter tuning and early stopping. All models were trained on the historical data prior to this validation window (January 1, 2020, to September 23, 2024).
Table 1: Temporal Partitioning of the Dataset.
Dataset Split Start Date End Date Length (Days)
Training January 1, 2020 September 23, 2024 1,728
Validation September 24, 2024 January 21, 2025 120
Test January 22, 2025 May 21, 2025 120
Total January 1, 2020 May 21, 2025 1,968
@@
The exogenous variable employed in this study is Global Liquidity, constructed as an aggregate measure of broad money supply across a diversified set of major and emerging economies. Specifically, the indicator includes M2 from the following economies, using the closest available national definition where necessary: the United States, Euro Area, China, Japan, United Kingdom, Canada, Australia, India, South Korea, Brazil, Russia, Switzerland, Mexico, Indonesia, Turkey, Saudi Arabia, Argentina, and South Africa.